# MineLand: Simulating Large-Scale Multi-Agent Interactions with Limited Multimodal Senses and Physical Needs

### v1

@@ Appendix A Broader Impact
Appendix B Details of MineLand Benchmark Suite
@@
The data statistics of the dataset are shown in Table B.1. Figure 6 displays specific task data.
Table B.1: Statistical analysis of the tasks in MineLand Benchmark Suite.
@@
Figure 6: Illustration of example for different tasks.
Appendix C Details of Alex
Different from most conventional LLM-based agents, Alex process information from various sources like visual, auditory (hearing conversations, environmental sounds), and tactile (touching objects) to build a comprehensive understanding of the world. For the remaining parts of MineLand that are not emphasized, we used the default setting of Wang et al. (2023). AI agents perceive the environment, make plans, execute plans independently, and interact with other agents. Internally, Alexmay exhibit different personality traits predefined in the system prompt. It utilizes the VLM for processing sensor information within the Minecraft environment. Alex also tracks its own states, energy levels, and resource inventory to inform its actions and prioritize tasks.
C.1 Memory Component
@@
The memory library is responsible for storing all of the agent’s information and retrieving relevant information from memory based on events. The memory library stores the agent’s personality, persona, long-term goals, short-term goals, chat records, experienced events, mastered skills, and environmental information. The memory library also includes a long-term planner. When accessing the memory library for the first time, the long-term planner generates a long-term plan based on the personality, persona, and observations, which is then stored in the memory library. In each iteration, the memory library processes information from observations, critic information, and tasks, and stores it in the vector database 777 We leverage Chroma in Alex for storage and retrieve memory. https://www.trychroma.com/. If the critic information indicates that the previous task has been completed, the skill manager in the memory library will be called upon to generate a concise description of the relevant skill, which will then be stored in the skill vector database.
@@
C.2 Hierarchical Planning Component
@@ C.3 Action Module
This self-correction circle allows for identifying and correcting deviations from planned behaviors or task execution. This equips the agent with mechanisms to detect and recover from errors like missed goals, failed actions, or unforeseen consequences. In previous work, such as Park et al. (2023), action execution will not fail. Planning to cook will definitely lead to success. However, in MineLand as well as reality, actions may fail due to various reasons, such as unexpected events. So, we need self-correction to solve some simple action errors in the action module. Inspired by Voyager, we implement self-correction. By reading information such as observation space and short-term plan, we comprehensively consider and determine the completion status of the short-term plan through the critic component.
C.4 Algorithm
Algorithm 1 demonstrates Alex’s core algorithm, which corresponds to the process in Figure 4.
Algorithm 1 Core algorithm of Alex.
  Input: observation space $obs$, task info $task$, code info $code$
  $state,\ description\leftarrow\text{Distributor}(obs,\ task,\ code)$
  if $state\ =\ None$ then
     return $None$
  end if
  if $state\ =\ Critic$ then
     $state,\ description,\ critic\leftarrow\text{Critic Component}(obs)$
  end if
  if $state\ =\ Action$ then
     $\text{Perceive}(obs,critic)$
  end if
  if $state\ =\ Brain$ then
     $\text{Perceive}(obs,critic,code)$
[…]
     $retrieved\leftarrow\text{Retrieve}(obs)$
     $\text{Plan}(obs,task,retrieved)$
     $state\leftarrow Action$
  end if
[…]
  return $act$
Appendix D Details of Hyper-Parameters
Alex leverage OpenAI’s gpt-4-vision-preview API for text completion in all components, and text-embedding-ada-002 API for text embedding in the memory library for storage and retrieve memory. Apart from the action component, the maximum tokens are set to 512. For the action component, the maximum tokens are expanded to 512 * 3.
@@ Appendix D Details of Hyper-Parameters
  • Dispatching component: “FAILED TIMES LIMIT” refers to the maximum number of attempts allowed when Alex has a code error. The default value is 5. “code execution time limit” is 2000 ticks.
  • Critic component: “FAILED TIMES LIMIT” refers to the maximum number of attempts allowed when Alex failed to achieve a short-term plan. The default value is 2. “Critic Mode” is “auto” for leveraging AI to judge whether to achieve the short-term plan.
  • Memory Library: “chat retrieve limit” is 5, “event retrieve limit” is 2, “environment retrieve limit” is 2, “skill retrieve limit” is 5, “recent chat retrieve limit” is 8, “short term plan retrieve limit” is 5.
Appendix E Minecraft Simulator Performance Comparison
We evaluate the number of agents that MineLand can support and compare MineLand with other popular Minecraft simulators, in Table A.1. We utilize a mainstream consumer desktop PC equipped with an Intel i5-12400F CPU and 64GB of memory. Vision condition means MineLand provides visual display and headless mode means MineLand doesn’t provide a visual display. The two values of the VISION column (e.g., 6 and 250MS) are view distance and visual refresh interval in milliseconds, respectively.
Our findings reveal that MineLand is capable of supporting up to 16 agents simultaneously while providing a visual display while it supports 48 agents without a visual display. Furthermore, when MineLand and Malmo both run 8 agents, MineLand’s CPU and memory usage are approximately 1/3 that of Malmo’s (specifically, 35.6% and 38.0%, respectively). It is worth noting that Malmo serves as the foundation for most popular Minecraft Platforms (e.g., MineDojo/MineRL v0.4/MarLÖ), thus highlighting MineLand’s superior performance compared to the vast majority of existing Minecraft Platforms.
It should be noted that original MineDojo and MineRL only support a single agent, while MarLÖ by default supports only 2 agents. Malmo originally supported only one agent. We have designed a new task for Malmo, so it supports 8 agents.
Table A.1: Comparison between MineLand and other popular Minecraft simulators under different conditions.
Simulator Conditions Number of Agents Avg CPU Time Max CPU Time Avg Mem Usage
@@ Appendix E Minecraft Simulator Performance Comparison
Appendix F Experiments of Single Agent
As shown in Figure 7, even with this restriction, Alex demonstrates a strong ability for long-term planning.
Appendix G Experiments of Physical Needs
We perform experiments using a survival task. The starting condition of the agent is a health level of 1 and a satiety level of 0, with an initial inventory of 2 bread. The objective for the agent is to survive for one day. Comparative experiments are carried out in regular MineLand and its counterpart version MineLand ${}_{w/o\ physical\ needs}$ without physical needs.
Alex in MineLand can survive one day in his shelter and successfully completed the task. We showcase an example of Alex, with short-term plans and actions generated chronologically by Alex:
@@ Appendix G Experiments of Physical Needs
  • Action: Mine 5 oak logs
@@
From the above plans and actions, we observe that, although Alex is aware of possessing bread, the lack of physical needs resulted in it not consuming the bread. Instead, the agent continuously collected resources to cope with threats. Due to the low health level, it was killed by a zombie at night, leading to the failure of the task.
@@
Repeated experiments on the constructed building task also show that Alex in MineLand with physical needs live longer. Agents tend to prioritize eating and then building a shelter, while Alex in MineLand without physical needs lean towards exploring and collecting resources. This indicates that physical needs’ importance to simulate real life.
Appendix H Experiments of Comparing VLMs and LLMs
@@
As shown in Table A.2, we perform thorough assessments using newly implemented LLMs/VLMs within our agent framework. The tasks used include “Harvest 1 White Wool With 1 Shears” and “Harvest 1 White Wool With 1 Iron Word”. We compared the performance of Alex, which uses the gpt-4-1106-preview API for the complete action component, and Alex, which uses different VLMs or LLMs for the action component.
@@
Table A.2: Comparison of different VLMs and LLMs for the action component. Fractions represent the count of successful completions within a set of three attempts. 0/3 means that the method is unable to solve the task within the maximum number of code iterations(15) or exceeds the designated area of the task. The fewer the number of code iterations, the higher the efficiency.
Task harvest 1 white wool with 1 shears harvest 1 white wool with 1 iron sword
@@
Appendix I Limitations

### v2

@@ Appendix B Broader Impact
Appendix C Details of MineLand Simulator
We show the comparison with related popular platforms or projects in Table 7.
Table 7: Comparison with related popular platforms or projects. Max Agents: The approximate maximum number of agents supported on a single PC (See Section 5.1 for details). Human: Whether the simulator allows humans to directly interact with AI agents. Plan as Action: Whether agents can generate plans in a specific format (e.g., code) that the simulator interprets and executes directly as actions. Sociological Experiments: Whether the simulator is designed to facilitate the study of social phenomena and emergent behavior in multi-agent systems. Physical Needs: Whether agents have simulated physiological needs (e.g., hunger, thirst) that influence their behavior and require actions to fulfill. Open World: Open-world or not. 3D Space: 3D environment or not.
Simulator Max Human Plan as Sociological Physical Open 3D Number
Agents Action Experiments Needs World Space of Tasks
MineLand 64+ 6000+
MineDojo [16] 1 - - - 3000+
MineRL v0.4 [20] 1 - - - 11
MineRL v1.0 [3] 1 - - - 5
MarLÖ [42] 8 - - 14
Malmo [27] 8 - - -
IGLU [30] 1 - - - 157
Voyager [57] 1 - -
Habitat 3.0 [47] 2+ - - - - 200
Habitat 2.0 [55] 1 - - - - - 105
VirtualHome-Social [46] 40+ - - 50
SAPIEN [62] 1 - - - - - -
AI2-THOR [31] 6 - - - - - 200+
iGibson [35] 2+ - - - - 50
Melting Pot [34] 50+ - - - - 256
Smallville [40] 25 - - -
C.1 Architecture
The architecture consists of three main modules: the Bot Module, the Environment Module, and the Bridge Module.
C.2 State Space
We focus on a state space that blends task-oriented activities with the rhythms of daily life. This blending opens up exciting possibilities for studying dynamic and ecologically valid multi-agent interactions. For example, agents have to balance the goal of mining with the need to fill their stomachs, otherwise they will starve to death. Here are detailed explanations for the state space:
C.3 Action Space
In contrast to generating textual action descriptions and training an external controller module [8] for transforming the plan to the executable code, this simulator is a language-model-friendly simulator, providing the code instead of the action description. Through the code, it can directly execute plans generated by the language model. Agents in MineLand generate a series of codes based on the Mineflayer API while representing planning, such as moving, watching, and mining. The advantage of using code is that it avoids error accumulation, whereas using a textual plan requires an additional model to map the plan to the code, which can lead to error accumulation.
C.4 Supporting up to 100 Agents
To evaluate the MineLand simulator’s ability to handle large-scale agents, we designed an experiment featuring 100 agents engaged in a simultaneous combat scenario (Figure 5). This high agent count pushes the boundaries of previously typical two-agent simulations and demonstrates the simulator’s scalability in effectively managing a massive number of agents acting concurrently.
Figure 5: 100 agents are fighting within the MineLand Simulator.
Appendix D Details of MineLand Benchmark Suite
@@
The data statistics of the dataset are shown in Table 8. Figure 6 displays specific task data. We show several blueprints in Figure 7.
Table 8: Statistical analysis of the tasks in the MineLand Benchmark Suite.
@@
Figure 6: Illustration of cases for different tasks.
Figure 7: Some of the blueprints for the construction tasks.
Appendix E Details of Alex
Crafting an AI agent for MineLand, where daily life seamlessly blends with task-oriented activities, opens up exciting possibilities. The agent should fulfill daily needs like cooking, socializing, and maintaining shelter, while also completing assigned tasks like resource gathering, crafting, or construction. We propose Alex, a VLM-based approach, to balance daily routines and tasks. Alex supports both individual daily-life goals and community-based task-oriented goals.
Our agent design addresses the unique requirements of our large-scale multi-agent MineLand simulator. Unlike prior simulators focused on smaller agent populations, our environment necessitates the incorporation of a social brain module. This module equips the agent with the ability to navigate complex and multitasking social interactions and make informed decisions within a crowded environment.
Furthermore, our Minecraft-inspired simulator introduces the challenge of action failure, unlike environments where “action spoken equals success” (referencing Smallville [40]). To address this, we have designed an action correction module. This module functions similarly to human neural reflexes, handling minor action corrections and alleviating the burden on the social brain module. This division of labor allows the social brain to focus on higher-level decision-making while the action correction module ensures smoother execution.
Different from most conventional LLM-based agents, Alex process information from various sources like visual, auditory (hearing conversations, environmental sounds), and tactile (touching objects) to build a comprehensive understanding of the world. For the remaining parts of MineLand that are not emphasized, we used the default setting of Voyager [57]. AI agents perceive the environment, make plans, execute plans independently, and interact with other agents. Internally, Alex may exhibit different personality traits predefined in the system prompt. It utilizes the VLM for processing sensor information within the Minecraft environment. Alex also tracks its own states, energy levels, and resource inventory to inform its actions and prioritize tasks.
E.1 Memory Component
@@
The memory library is responsible for storing all of the agent’s information and retrieving relevant information from memory based on events. The memory library stores the agent’s personality, persona, long-term goals, short-term goals, chat records, experienced events, mastered skills, and environmental information. The memory library also includes a long-term planner. When accessing the memory library for the first time, the long-term planner generates a long-term plan based on the personality, persona, and observations, which is then stored in the memory library. In each iteration, the memory library processes information from observations, critic information, and tasks, and stores it in the vector database 999 We leverage Chroma in Alex for storing and retrieving memory. https://www.trychroma.com/. If the critic information indicates that the previous task has been completed, the skill manager in the memory library will be called upon to generate a concise description of the relevant skill, which will then be stored in the skill vector database.
@@
E.2 Hierarchical Planning Component
@@ E.3 Action Module
This self-correction circle allows for identifying and correcting deviations from planned behaviors or task execution. This equips the agent with mechanisms to detect and recover from errors like missed goals, failed actions, or unforeseen consequences. In previous work, such as Smallville [40], action execution will not fail. Planning to cook will definitely lead to success. However, in MineLand as well as reality, actions may fail due to various reasons, such as unexpected events. So, we need self-correction to solve some simple action errors in the action module. Inspired by Voyager [57], we implement self-correction. By reading information such as observation space and short-term plan, we comprehensively consider and determine the completion status of the short-term plan through the critic component.
E.4 Multitasking Mechanism
We’ll explain the multitasking mechanism from both the simulator side and the agent side. From the simulator-side event management, the simulator maintains a dedicated event queue that stores various special events categorized by their urgency (e.g., hurt events, chat events, and death events). In terms of agents, when an agent encounters an event from the queue, it employs a prioritization mechanism to assess its importance and urgency. Lower-priority events, such as chat events, can be deferred to minimize disruption of the ongoing task. The agent’s decision-making module (referred to as the brain module) plays a central role. If the event necessitates attention, the brain module can either:
  • Context Switching: The brain module temporarily stores the ongoing task in the memory library and focuses on addressing the new event. This facilitates context switching when necessary.
  • Concurrent Processing: In specific scenarios, the brain module might choose to handle both tasks simultaneously, leveraging its multitasking capabilities.
This strategy, involving the simulator’s event queue and the agent’s prioritization and execution mechanisms, fosters effective multitasking within the multi-agent environment.
Appendix F Details of Hyper-Parameters
Alex leverage OpenAI’s gpt-4-vision-preview[39] API with a temperature of 0 for text completion in all components, and text-embedding-ada-002 API for text embedding in the memory library for storage and retrieve memory. Apart from the action component, the maximum tokens are set to 512. For the action component, the maximum tokens are expanded to 512$\times$3.
@@ Appendix F Details of Hyper-Parameters
  • Dispatching component: “FAILED TIMES LIMIT” refers to the maximum number of attempts allowed when Alex has a code error. The default value is 3. “code execution time limit” is 2000 ticks.
  • Critic component: “FAILED TIMES LIMIT” refers to the maximum number of attempts allowed when Alex failed to achieve a short-term plan. The default value is 2. “Critic Mode” is “auto” when we allow agents to automatically judge whether to achieve the short-term plan.
  • Memory Library: “chat retrieve limit” is 5. “event retrieve limit” is 2. “environment retrieve limit” is 2. “skill retrieve limit” is 5. “recent chat retrieve limit” is 8. “short term plan retrieve limit” is 5.
  • MineLand & Minecraft Server: By default, “gamemode” is survival mode. “difficulty” is peaceful mode (normal in combat tasks). “view-distance” on server side is 6. “simulation-distance” is 3. “task mode” is cooperative mode.
Appendix G Experiments of Minecraft Simulator Performance Comparison
We evaluate the number of agents that MineLand can support and compare MineLand with other popular Minecraft simulators, in Table 9. We utilize a mainstream consumer desktop PC equipped with an Intel i5-12400F CPU and 64GB of memory. Vision condition means MineLand provides visual display and headless mode. The headless mode means MineLand doesn’t provide a visual display. The two values of the Vision column (e.g., 6 and 250ms) are view distance and visual refresh interval in milliseconds, respectively.
Our findings reveal that MineLand is capable of supporting up to 32 agents simultaneously while providing a visual display, and up to 128 agents in the headless mode without visual displays. Furthermore, when MineLand and Malmo both run 8 agents, MineLand’s CPU and memory usage is approximately 1/3 that of Malmo’s (specifically, 35.6% and 38.0%, respectively). It is worth noting that Malmo serves as the foundation for most popular Minecraft Platforms (e.g., MineDojo/MineRL v0.4/MarLÖ), thus highlighting MineLand’s superior performance compared to the vast majority of existing Minecraft Platforms.
It should be noted that original MineDojo and MineRL only support a single agent, while MarLÖ by default supports only 2 agents. Malmo originally supported only one agent. We have designed a new task for Malmo to test the maximum number of agents it can handle. The result is 8.
Table 9: Comparison of performance between MineLand and other popular Minecraft simulators under different conditions. MineLand* represents an optimized version, so its Max CPU is significantly reduced.
Simulator Conditions Agents Avg CPU Max CPU Avg Mem
@@ Appendix G Experiments of Minecraft Simulator Performance Comparison
MineLand* Headless Mode 64 2.87% 18.84% 6.30 GB
MineLand* Headless Mode 128 4.00% 26.95% 9.04 GB
@@
Appendix H Experiments of Multimodal Observation
We leverage OpenAI’s gpt-4-vision-preview API for text completion and text-embedding-ada-002 API for text embedding. The temperature is set to 0. Unless specified, all experiments in Section 5 are set to this default setting. See Section F for details.
[…]
Simulator Time Success Rate
MineLand 46.38s 80%
[…]
Appendix I Experiments of Reinforcement Learning Agent
While our MineLand simulator is primarily designed for agents powered by LLMs or VLMs, it also offers functionalities that seamlessly integrate with Reinforcement Learning (RL) modules. This enables the training and evaluation of RL agents within the simulator’s environment. We conducted experiments of an example RL agent to show this functionality. We developed a RL agent using MineLand and Stable Baselines3 [51]. We fed the encapsulated gym-style MineLand directly into Stable Baseline3 and train the RL agent in low-level action mode. For other settings of Stable Baseline3, we follow the default setting. We adopt the “combat 1 zombie” task and the amount of blood volume lost by the zombie is award. The agent is trained using RL, as illustrated by the provided reward curve (Figure 9). The upward trajectory of the reward curve indicates that the RL agent is effectively learning strategies that result in higher rewards. The curve’s stabilization in the later stages of training suggests that the RL agent has converged to a stable policy, which demonstrates its superiority over the untrained agent.
Again, our core focus lies in the LLM/VLM domain. The simulator offers functionalities that can accommodate RL modules as an ancillary feature. This expands the potential applications of the simulator by enabling the training and evaluation of RL agents within its environment, albeit as a secondary capability.
Figure 9: Illustration of the reward curve in “combat 1 zombie” task, using DQN with MLP policy, $3000$-$5000$ total time steps.
Appendix J Experiments of Limited Senses
To assess the impact of limited senses, we carry out the “shearing the sheep” task. An agent with a first perspective (shown in Figure 10) is required to locate a sheep to complete the task. In the case of an agent with limited visual senses, we set 6 as the view distance and 250ms as the visual refresh interval in milliseconds. If there are no sheep within its field of view, it cannot shear a sheep. However, the agent can communicate with other agents to find a sheep. On the contrary, agents with unlimited visual senses will see distant sheep, prompting them to catch that sheep. Note that the unlimited visual senses here refer to a long view distance. If there is an obstruction, the agent still cannot see what is behind the obstruction.
Figure 10: Visual observation of agents with a first perspective.
Appendix K Experiments of Physical Needs
We perform experiments using a survival task. The starting condition of the agent is a health level of 1 and a satiety level of 0, with an initial inventory of 2 breads. The objective for the agent is to survive for one day. Comparative experiments are carried out in regular MineLand and its counterpart version MineLand ${}_{w/o\ physical\ needs}$. Alex in MineLand can survive one day in his shelter and successfully complete the task while the same agent in MineLand ${}_{w/o\ physical\ needs}$ failed to execute the task. The failure is due to the lack of physical needs, although Alex is aware of possessing bread, the lack of physical needs resulted in it not consuming the bread. Instead, the agent continuously collected resources to cope with threats. Due to the low health level, it was killed by a zombie at night, leading to the failure of the task. We showcase the details as follows:
We showcase an example of Alex, with short-term plans and actions generated chronologically by Alex:
@@ Appendix I Experiments of Reinforcement Learning Agent
  • Action: Mine 5 oak logs.
@@
Repeated experiments on the constructed building task also show that Alex in MineLand with physical needs live longer. Agents tend to prioritize eating and then building a shelter, while Alex in MineLand ${}_{w/o\ physical\ needs}$ towards exploring and collecting resources. This indicates that physical needs are important to simulate real life.
Appendix L Experiments of Single Agent
We assess individual Alex’s capabilities based on tech tree tasks, demonstrating that Alex’s architecture can plan and execute complex tasks. We attempt the task of obtaining diamonds six times, unlocking crucial items such as the crafting table, wooden pickaxe, stone pickaxe, iron ore, coal, furnace, iron ingot, and iron pickaxe in the process. Alex get diamonds twice out of six tasks. Importantly, to approach the real world, we’ve added a restriction on multimodal information, where Alex can obtain the location of the target only when it can visually discover or reason about the target’s presence nearby. This is because our agent Alex has the feature of limited senses. However, Voyager can “cheat”. Voyager obtains the location of the target directly from the system even though they have not seen the target.
As shown in Figure 11, even with this restriction, Alex demonstrates a strong ability for long-term planning. Additionally, when a plan does not yield the desired results (e.g., suddenly encountering obstacles on the road), Alex promptly adjusts and devises a short-term plan using its multimodal information.
Appendix M Experiments of Construction Tasks
We show the results of our un-finetuned Alex in two example construction tasks: “Monument Construction” and “Stone Stele Construction” (detailed in Figure 12). While Alex demonstrates a promising capability in selecting appropriate materials for both tasks, its overall construction abilities are currently limited.
Figure 12: Illustration of Alex completing the construction task.
Appendix N Experiments of Stage Performance Tasks
We show the results of stage performance tasks in Table 11, Table 12, and Table 13. We omit the results of “Romeo and Julia, Act I Scene I” stage performance task. Due to the large number of characters appearing (up to 13), our Alex  cannot handle such a complex performance task, resulting in poor performance results and a lack of readable action sequence.
Table 11: Illustration of Alex completing the “Make friends” stage performance task.
Script Name Make friends
Agent number 3
Script There are three agents: Bob, Alice, and Jack. They want to make friends. Bob says, ‘Hello.’ Alice responds, ‘Hello.’ Jack says, ‘It’s good weather; let’s make friends!’ Bob agrees, and Alice says, ‘Nice.’
$SEQ^{*}$ Bob chat ‘Hello’. Alice chat ‘Hello. Jack chat ‘It’s good weather, lets make friends!’. Bob chat ‘I agree.’. Alice chat ‘nice’.
$SEQ_{Agent}$ Bob chat ‘Hello’. Alice chat ‘Hello Bob!’. Jack chat ‘Hello Bob!’. Jack chat ‘It’s good weather, let’s make a friend!’. Bob chat ‘I agree.’. Jack chat ‘Bob has agreed!’. Alice chat ‘nice’.
Keypoing Score 0.98
Appropriateness Score 0.67
Human Evaluation Score 3
Reason Agents’ words are too redundant, which made the performance a little stiff.
Table 12: Illustration of Alex completing the “Exchange items” stage performance task.
Script Name Make friends
Agent number 2
Script Bob has shears, and Alice has an iron sword. Bob says, ‘I want to exchange my shears for your iron sword.’ Alice agrees. Bob gives Alice the shears first, and then Alice gives the iron sword to Bob.
$SEQ^{*}$ Bob chat ‘I want to exchange my scissors for your iron sword’. Alice chat ‘I agree’. Alice get shears. Bob get $iron\_sword$.
$SEQ_{Agent}$ Bob chat ‘Hello Alice, I am here for the exchange’. Bob chat ‘I want to exchange my scissors for your iron sword’. Alice chat ‘Hello Bob, I am ready for the exchange.’. Alice get shears. Bob get $iron\_sword$.
Keypoing Score 0.99
Appropriateness Score 0.59
Human Evaluation Score 4
Reason The agents said something unnecessary.
Table 13: Illustration of Alex completing the “Cook food” stage performance task.
Script Name Exchange items
Agent number 1
Script Bob is hungry, and now there is a pig in the presence of of him, he want to kill the pig, get the porkchop, cook and eat it.
$SEQ^{*}$ pig died. Bob get porkchop. Bob get cooked_porkchop. Bob eat cooked_porkchop.
$SEQ_{Agent}$ pig died. Bob get porkchop. Bob get cooked_porkchop. Bob eat cooked_porkchop.
Keypoing Score 1.00
Appropriateness Score 1.00
Human Evaluation Score 5
Reason The agent’s behavior is completely consistent with the script.
Appendix O Experiments of Simulating Sociological Phenomena
O.1 Experiments of Simulating Conformity Phenomena
We utilize the task “Choose the Taller Tower” to demonstrate that MineLand is capable of supporting sociological experiments with more than 10 agents. In this task, we placed multiple agents (1, 2, 5, 10 agents) in a line in front of two towers of different heights and asked them to choose the taller one. All agents except the last one (the agent to test) deliberately chose the shorter tower. After observing the actions of the other agents, the last agent made its own choice. The last agent could be set as two types of personalities: one that tends to conform to others and one that tends to make independent judgments. As shown in Table 14, the last agent with a conformist personality chose the answer chosen by the majority in all scenarios (even though it is a wrong answer) except when the agent is the only one doing this task. Participants with an independent personality always chose the correct answer, regardless of others’ opinions.
Through this experiment, we demonstrated that MineLand can simulate sociological experiments with large-scale agents and accurately exhibit behaviors corresponding to different personalities.
Table 14: Results of “Choose the Taller Tower”. 4+1 means there are 5 agents in the MineLand, the former 4 of them are confederates and the last 1 is the agent to test. Choosing the “right” tower means the last agent chooses the correct tower.
Agent Type Number of Agents
0+1 1+1 4+1 9+1
Conformist Agent right left left left
Non-conformist Agent right right right right
O.2 Experiments of Simulating Personality Traits
We utilize the task “Treasure Hunt in the Forest” to show that MineLand can support sociological experiments involving 48 agents simultaneously. In the task “Treasure Hunt in the Forest”, agents with different levels of extraversion and agreeableness personality traits planned their actions. We used the long-term plans generated by the agents to assess whether they exhibited a tendency to cooperate with others. As shown in Table 15, all agents in the high-level group showed a tendency to cooperate, explicitly stating in their long-term plans that they intended to work with others. In the low-level group, only a small portion of agents exhibited a tendency to cooperate. Although the task information indicated that it was a cooperative task (prompts with the task information “Interact with other players if necessary” and “Minimize interactions with other players.”), many agents expressed reluctance to cooperate. This result is consistent with the psychology theory: five-factor model of personality [13].
The results also demonstrate that our MineLand can support up to 48 agents and that the combination of agents and the MineLand can simulate social phenomena and conduct meaningful sociological experiments.
Table 15: Cooperative tendency in agents with different levels of extraversion and agreeableness. We show the number of agents who exhibited a tendency to cooperate with others.
Agent Type Number of Cooperative Agents
High-level 48
Low-level 24
Appendix P Experiments of Comparing Different VLMs and LLMs
@@
As shown in Table 16, we perform thorough assessments using newly implemented LLMs/VLMs within our agent framework. The tasks used include “Harvest 1 White Wool With 1 Shears” and “Harvest 1 White Wool With 1 Iron Word”. We compared the performance of Alex, which uses different VLMs or LLMs for the action component.
@@
Table 16: Comparison of different VLMs and LLMs for the action component. Fractions represent the count of successful completions within a set of three attempts. 0/3 means that the method is unable to solve the task within the maximum number of code iterations (15) or exceeds the designated area of the task. The fewer the number of code iterations, the higher the efficiency.
LLM or VLM Task
Harvest 1 white wool with 1 shears Harvest 1 white wool with 1 iron sword
@@
Appendix Q Evaluation Metrics For Construction Tasks
For construction tasks, we introduced VLM-based and human evaluation methods to score the agent. The score is an integer ranging from 1 to 5, with higher scores indicating that the agents’ constructions more closely match the blueprint. The scoring criteria for construction tasks are shown in Table 17. We show examples of evaluation conducted by a GPT-4 model as a judge for the “Construct A Stone Stele” task (Figure 14) and the “Construct A Monument” task (Figure 15). We display the evaluation prompt in Figure 13.
Table 17: Scoring criteria for construction tasks.
Score Evaluation Criteria
5 The construction perfectly matches the blueprint, with all elements accurately placed and fully completed.
4 The construction mostly matches the blueprint, with minor inaccuracies and nearly complete.
3 The construction somewhat matches the blueprint, with noticeable inaccuracies and partially complete.
2 The construction partially matches the blueprint, with significant inaccuracies and largely incomplete.
1 The construction does not match the blueprint, with major inaccuracies and mostly incomplete.
Figure 13: Prompt for VLM-based evaluation.
Figure 14: An example of evaluation conducted by a GPT-4 model as a judge for the “Construct A Stone Stele” task.
Figure 15: An example of evaluation conducted by a GPT-4 model as a judge for the “Construct A Monument” task.
Appendix R Evaluation Metrics For Stage Performance Tasks
For stage performance tasks, we introduced human evaluation methods to score the agent. The score is an integer ranging from 1 to 5, with higher scores indicating that the agent’s performance is more aligned with the script and more natural. The scoring criteria for stage performance tasks are shown in Table  18.
Table 18: Scoring criteria for stage performance tasks.
Score Evaluation Criteria
5 The action of agent(s) completely matches the script, the performance is smooth and natural, emotions are expressed accurately, interactions with other characters or the environment are natural.
4 The action of agent(s) mostly matches the script, the performance is generally smooth with few inconsistencies, emotions are expressed accurately, interactions are natural.
3 The action of agent(s) roughly matches the script, the performance has some continuity but noticeable pauses, emotions are mostly accurate, interactions are somewhat stiff.
2 The action of agent(s) partially matches the script, the performance lacks continuity with many pauses, emotions are not expressed accurately, interactions are unnatural.
1 The action of agent(s) does not match the script, the performance is very disjointed, emotions are not expressed accurately, interactions are stiff.
Appendix S Limitations